# EF-3DGS: Event-Aided Free-Trajectory 3D Gaussian Splatting

### v3

@@ Abstract
Figure 1: Free-trajectory 3DGS under high speed. The top row shows the overall paradigm. The colored dots (red for positive and blue for negative events) in the top row represent the event data. We leverage continuous event streams to aid discrete video frames captured along free trajectories in high-speed scenarios, jointly optimizing camera poses and reconstructing the 3DGS. Our method surpasses current state-of-the-art methods in terms of both rendered results (middle row) and pose estimation (bottom row).
@@ 1 Introduction
In recent years, Neural Radiance Fields (NeRF) [30, 1, 31] and 3D Gaussian splatting (3DGS) [20, 9] have made significant progress in novel view synthesis tasks. Given a set of posed images of the same scene, they optimize an implicit or explicit scene representation using volume rendering. While subsequent methods [1, 31, 7, 2] excel with posed images, reconstructing scenes from videos with free camera trajectories remains challenging despite its applications in VR/AR, video stabilization, and mapping. To tackle this challenging task, several efforts have been made.
@@
Accurate pose estimation is often difficult to obtain in free-trajectory scenarios, which directly impacts the quality of scene reconstruction. One line of work draws inspiration from Simultaneous Localization and Mapping (SLAM). They [28, 9, 19] follow its optimization paradigm, progressively optimizing camera trajectories and alternating between camera pose and scene refinement. Another line of work [4, 19, 25, 9, 42] explores incorporating additional geometric or motion priors such as depth estimation [29, 36] or optical flow [41] to establish constraints beyond photometric rendering loss. While these methods can render photo-realistic images in typical free-trajectory scenarios, both their rendering quality and pose estimation accuracy degrade significantly in high-speed scenarios (or equivalently low-frame-rate scenarios) as shown in Fig. 1. Such high-speed scenarios have essential applications such as autonomous driving and First-Person View (FPV) exploration.
@@
The performance degradation of prior methods can be attributed to two primary factors. First, the limited number of camera observations leads to an under-constrained scene reconstruction problem. This can cause the scene representation to converge to a trivial solution [32, 47, 42], where the model overfits to the training views without capturing the correct underlying geometry structure. Second, the substantial discrepancies between consecutive frames, resulting in diminished overlapping regions, violate the implicit assumption of continuous motion between adjacent frames, which is leveraged by previous methods. Moreover, geometric and motion priors like optical flow and feature matching become unreliable in such scenarios. These significant violations greatly exacerbate the ill-posedness of the joint optimization of scene and camera poses.
@@
Event camera is a bio-inspired image sensor that asynchronously records per-pixel brightness changes, offering advantages such as high temporal resolution, high dynamic range, and no motion blur [12, 17, 45, 46, 52, 6]. The brightness information recorded in the event stream can effectively complement the missing scene information between consecutive frames. Moreover, the event data naturally encodes the motion information of the scene [24, 40, 44], containing rich motion cues. These properties make event cameras well-suited for scene reconstruction tasks in high-speed and free-trajectory scenarios. However, seamlessly integrating the aforementioned benefits of event cameras is nontrivial. First, 3DGS renders absolute pixel brightness, which aligns with image data. Event cameras, however, record sparse differential brightness changes. Directly integrating the differential operations into 3DGS may amplify noise and lead to ill-conditioned optimization problems with high sensitivity to parameter initialization and perturbations. Second, event cameras encode motion through continuous spatio-temporal trajectories of events. In contrast, frame-based data inherently discretizes continuous motion, forcing traditional methods to rely on correspondence matching, which fails in high-speed scenarios with large inter-frame displacements. These fundamental challenges require carefully designed method that bridges the gap between the event data and 3DGS optimization.
@@
In this work, we propose Event-Aided Free-Trajectory 3DGS, dubbed EF-3DGS, a framework that integrates event data into the scene optimization process to fully leverage its high temporal resolution property. Our approach comprises three key components: (1) In the Event Generation Model (EGM), we introduce an event-based re-render loss, which extends the 3DGS optimization to the continuous event stream. This allows us to utilize the brightness cues encoded in the event stream between adjacent frames, providing rich supervisory signals to alleviate the insufficient sparse view issues. (2) In the Linear Event Generation Model (LEGM), regarding the pose estimation challenge, we introduce the CMax [11] framework to exploit the spatio-temporal correlations of events. We obtain the motion field by leveraging the pseudo-depth from 3DGS rendering and the relative camera motion between consecutive frames. We then warp the events triggered by the same edge along the motion trajectories to maximize the sharpness of the image of warped events (IWE), thereby estimating the motion that best matches the current spatio-temporal event patterns. Furthermore, through the Linear Event Generation Model [10, 13], we establish a connection between motion and brightness changes. This allows us to constrain the 3DGS in the gradient domain using the IWE. (3) As most event data primarily records scene brightness changes, lacking color information, we introduce photometric bundle adjustment (PBA) and a Fixed-GS strategy to address this. PBA recovers color by optimizing reprojection errors onto RGB frames, while Fixed-GS enables separate optimization of scene structure and color.
@@ 2.1 Joint Pose and Scene Optimization
The research community has recently focused on developing methods [42, 28, 50, 8, 4, 23, 9] that can be optimized without requiring precomputed camera poses. A line of work has focused on improving the stability of the optimization process. GARF [8] and BARF [23] both find that the high-frequency position encoding is prone to local minima and try to improve it. For example, GARF [8] proposes using Gaussian activation to replace the sinusoidal position encoding. Another line of work has investigated incorporating additional constraints to make the problem more tractable. LocalRF [28] leverages the prior assumption of continuous motion between adjacent frames and progressively adds and optimizes camera poses. More recent approaches [4, 28, 9] leverage pre-trained networks, i.e., monocular depth estimation and optical flow estimation. Exploiting 3DGS’s explicit representation, CF-3DGS [9] directly back-projects Gaussian points using depth maps. While the aforementioned methods have made notable progress, they have yet to fully address the challenges posed by high-speed scenarios or rely on a good pose initialization. Our approach addresses these issues by leveraging motion and brightness cues from event streams.
2.2 Event-Based Novel View Synthesis
@@
Recent works have explored the integration of event cameras into the NeRF or 3DGS framework. Early approaches, such as E-NeRF [21] and EventNeRF [37], utilize event-based generative models, minimizing the difference between the rendered brightness changes and observed brightness changes. Building upon this, Robust e-NeRF [26] incorporates a more realistic imaging model into the event-based framework, accounting for factors like refractory periods and noise. Beyond event-based NeRF, efforts have also been made to integrate event data into image-based methods. For instance, E2NeRF [35] and EvDeblurNeRF [5] leverage the Event Double Integral (EDI) [33] model to address the deblurring problem, while DE-NeRF [27] and EvDNeRF [3] leverage the high temporal resolution property of event cameras to capture fast-moving elements in dynamic scene. More recently, Event-3DGS [16] and EaDeblur-GS [51] have extended previous approaches to 3D Gaussian Splatting, achieving superior rendering quality and real-time performance. A key distinction of our work is that, unlike the prior methods that rely on accurate precomputed poses, we target free-trajectory scenarios, jointly optimizing for both the camera poses and the scene representation. Furthermore, while previous works have been limited to simulated and simple environments, we evaluate our approach in large-scale outdoor scenarios with complex motions and lighting conditions.
@@
Figure 2: Method overview. Our method takes video frames and event stream as input. During the first stage, we progressively add new event images, leveraging the events and most recent frame to establish the event-driven optimization. During the second stage, we adopt the Fixed-GS strategy to mitigate the color distortion of 3DGS. The details of $\mathcal{L}_{LEGM}$ and CMax framework are illustrated in Fig. 3.
@@ 3 Preliminary
3DGS [20] parametrizes the 3D scene as a set of 3D gaussians $\{G_{k}\}_{k=1}^{K}$ that carry the geometric and appearance information. Each 3D Gaussian is characterized by several learnable properties, including its center position $\mu\in\mathbb{R}^{3}$, opacity $\alpha\in[0,1]$, spherical harmonics (SH) features $\mathbf{f}_{k}\in\mathbb{R}^{3\times 16}$ for view-dependent color $c\in\mathbb{R}^{3}$, rotation matrix $R\in\mathbb{R}^{3\times 3}$ (stored in quaternion form), scale factor $s\in\mathbb{R}^{3}$. The shape of each Gaussian is defined by the covariance matrix $\mathbf{\Sigma}$ and the center (mean) point $\mu$, $G(x)=\exp(-\frac{1}{2}{(x-\mu)}^{T}\mathbf{\Sigma}^{-1}(x-\mu)).$ During rendering, a tile-based rasterizer is applied to enable fast sorting and $\alpha$-blending. The color of each pixel is calculated via blending N ordered overlapping points:
@@
where $c_{i}$ is calculated from spherical harmonics and view direction, $\alpha_{i}$ is the multiplication of opacity and the transformed 2D Gaussian and r denotes the image pixel. With the forward rendering procedure, we can optimize 3DGS by minimizing a weighted combination loss of $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ between observation and rendered pixels: $\mathcal{L}_{color}=(1-\lambda)\mathcal{L}_{1}(\hat{I},I)+\lambda\mathcal{L}_{D-SSIM}(\hat{I},I),$ where $\lambda$ is balancing weight which is set to 0.2 following [20]. By integrating depth $d_{i}$ in Equation 1 along the ray, we can also obtain a expected depth value $\hat{D}(\mathbf{r})$:
@@ 4 Method
The overall framework is shown in Fig. 2. Given a video of a free-trajectory $\{I_{i}\}$ captured at time $\{t_{i}\}$ and the event stream $\varepsilon=\{\mathbf{e}_{k}\}$, our goal is to reconstruct the 3DGS of the scene and the corresponding camera trajectory$\{T_{i}\}$. Following the analysis-by-synthesis paradigm of 3DGS, we extend this approach by incorporating event camera data through two fundamental imaging principles: the Event Generative Model (EGM) and Linear Event Generative Model (LEGM). To address the absence of color information in events and ensure cross-view consistency, we further introduce photometric bundle adjustment (PBA) and a Fixed-GS training strategy.
@@ 4.1 EGM Driven Optimization
The Event Generation Model (EGM) describes how event cameras asynchronously record pixel-wise brightness changes. When the logarithmic brightness change at a pixel $\mathbf{u}_{k}=(x_{k},y_{k})$, exceeds a predefined contrast threshold $C$,
@@ 4.2 Unified CMax and LEGM Optimization
While $\mathcal{L}_{EGM}$ leverages the brightness change information recorded by events, it does not explicitly exploit the motion information encoded in the event stream. To address this, we introduce the Contrast Maximization (CMax) [11, 34, 38, 15] framework and the Linear Event Generation Model (LEGM) [10, 13, 56]. These models complement the previous EGM-driven optimization.
Under constant scene illumination, events are triggered by the motion of scene edges, forming continuous trajectories in (x, y, t) space. As shown in Fig. 2 (1.2), by warping(back-projecting) events along the correct motion trajectories, we can obtain a sharp image of warped events (IWE). Therefore, the sharpness of the IWE can serve as an indication of the accuracy of the estimated motion. This insight motivates us to derive the motion field by leveraging the rendered depth from 3DGS using Eq. 2 and the relative camera motion between neighboring timestamps. By optimizing the sharpness of the IWE, we can obtain the optimal motion field, which in turn helps to improve the geometric accuracy of the 3DGS and the camera poses.
@@
$\vspace{1pt}E_{i,j-m\rightarrow j}=\textrm{warp}(E_{i,j-m},F_{i,j\rightarrow j-m}),\vspace{1pt}$ (7)
where $m\in[0,r]$, $F_{i,j-m\rightarrow j}$ is the optical flow derived from the rendered depth $\hat{D}$ in  Eq. 2 and relative pose $T_{i,j\rightarrow j-m}$ between two timestamps:
$\vspace{1pt}F_{i,j\rightarrow j-m}=\Pi(T_{i,j\rightarrow j-m}\Pi^{-1}(x,y,\hat{D}))-(x,y),\vspace{1pt}$ (8)
@@
Furthermore, based on the Linear Event Generation Model (LEGM) [10, 13], the brightness change $\Delta L$ at pixel $\mathbf{u}$ can be approximated by the dot product of the image gradient $\nabla L$ and the optical flow $\dot{\mathbf{u}}$ (note that $L$ is the logarithm of an image):
@@
It is noteworthy that the IPWE also encodes brightness change information. Combining  Eq. 9 and  Eq. 11, we establish a connection between the IPWE and the brightness changes of the rendered images:
@@
Note that to compute $F_{i,j\rightarrow j+1}$, we estimate $T_{i,j,j+1}$ by leveraging the assumption of locally linear motion from $T_{i,j,j-1}$ and $T_{i,j,j}$. Based on this relationship, we formulate an additional gradient-based loss:
$\vspace{1pt}\mathcal{L}_{grad}=||C\cdot\textrm{IPWE}_{i,j}-(\hat{L}(\mathbf{u})-\hat{L}(u+F_{i,j\rightarrow j+1}))||^{2},\vspace{1pt}$ (13)

### v4

@@ Abstract
Figure 1: Free-trajectory 3DGS under high speed. (Top) The overall paradigm. The colored dots in the top row represent the event data (red: positive, blue: negative). We leverage continuous event streams to aid discrete video frames captured along free trajectories in high-speed scenarios, jointly optimizing camera poses and reconstructing the 3DGS. Our method surpasses current state-of-the-art methods in terms of both rendered results (middle) and pose estimation (bottom).
@@ 1 Introduction
In recent years, Neural Radiance Fields (NeRF) nerf ; mipnerf360 ; instantngp and 3D Gaussian splatting (3DGS) 3dgs ; cf3dgs have made significant progress in novel view synthesis tasks. Given a set of posed images of the same scene, they optimize an implicit or explicit scene representation using volume rendering. While subsequent methods mipnerf360 ; instantngp ; tensorf ; zipnerf excel with posed images, reconstructing scenes from videos with free camera trajectories remains challenging despite its applications in VR/AR, video stabilization, and mapping. To tackle this challenging task, several efforts have been made.
@@
Accurate pose estimation is often difficult to obtain in free-trajectory scenarios, which directly impacts the quality of scene reconstruction. One line of work draws inspiration from Simultaneous Localization and Mapping (SLAM). They localrf ; cf3dgs ; cogs follow its optimization paradigm, progressively optimizing camera trajectories and alternating between camera pose and scene refinement. Another line of work nopenerf ; cogs ; rodynrf ; cf3dgs ; sparf explores incorporating additional geometric or motion priors such as depth estimation depth1 ; depth2 or optical flow raft to establish constraints beyond photometric rendering loss. While these methods can render photo-realistic images in typical free-trajectory scenarios, both their rendering quality and pose estimation accuracy degrade significantly in high-speed scenarios (or equivalently low-frame-rate scenarios) as shown in Fig. 1. Such high-speed scenarios have essential applications such as autonomous driving and First-Person View (FPV) exploration.
@@
The performance degradation of prior methods can be attributed to two primary factors. First, the limited number of camera observations leads to an under-constrained scene reconstruction problem. This can cause the scene representation to converge to a trivial solution regnerf ; sparsenerf ; sparf , where the model overfits to the training views without capturing the correct underlying geometry structure. Second, the substantial discrepancies between consecutive frames, resulting in diminished overlapping regions, violate the implicit assumption of continuous motion between adjacent frames, which is leveraged by previous methods. Moreover, geometric and motion priors like optical flow and feature matching become unreliable in such scenarios. These significant violations greatly exacerbate the ill-posedness of the joint optimization of scene and camera poses.
@@
Event camera is a bio-inspired image sensor that asynchronously records per-pixel brightness changes, offering advantages such as high temporal resolution, high dynamic range, and no motion blur event_survey ; EDSO ; ultimateslam ; wan2024event ; wu2024event ; chengzhi . The brightness information recorded in the event stream can effectively complement the missing scene information between consecutive frames. Moreover, the event data naturally encodes the motion information of the scene EvMotionEstimation ; ganchao ; 10047993 , containing rich motion cues. These properties make event cameras well-suited for scene reconstruction tasks in high-speed and free-trajectory scenarios. However, seamlessly integrating the aforementioned benefits of event cameras is nontrivial. First, 3DGS renders absolute pixel brightness, which aligns with image data. Event cameras, however, record sparse differential brightness changes. Directly integrating the differential operations into 3DGS may amplify noise and lead to ill-conditioned optimization problems with high sensitivity to parameter initialization and perturbations. Second, event cameras encode motion through continuous spatio-temporal trajectories of events. In contrast, frame-based data inherently discretizes continuous motion, forcing traditional methods to rely on correspondence matching, which fails in high-speed scenarios with large inter-frame displacements. These fundamental challenges require carefully designed method that bridges the gap between the event data and 3DGS optimization.
@@
In this work, we propose Event-Aided Free-Trajectory 3DGS, dubbed EF-3DGS, a framework that integrates event data into the scene optimization process to fully leverage its high temporal resolution property. Our approach comprises three key components: (1) In the Event Generation Model (EGM), we introduce an event-based re-render loss, which extends the 3DGS optimization to the continuous event stream. This allows us to utilize the brightness cues encoded in the event stream between adjacent frames, providing rich supervisory signals to alleviate the insufficient sparse view issues. (2) In the Linear Event Generation Model (LEGM), regarding the pose estimation challenge, we introduce the CMax cmax framework to exploit the spatio-temporal correlations of events. We obtain the motion field by leveraging the pseudo-depth from 3DGS rendering and the relative camera motion between consecutive frames. We then warp the events triggered by the same edge along the motion trajectories to maximize the sharpness of the image of warped events (IWE), thereby estimating the motion that best matches the current spatio-temporal event patterns. Furthermore, through the LEGM legm2015 ; legm2018 , we establish a connection between motion and brightness changes. This allows us to constrain the 3DGS in the gradient domain using the IWE. (3) As most event data primarily records scene brightness changes, lacking color information, we introduce photometric bundle adjustment (PBA) and a Fixed-GS strategy to address this. PBA recovers color by optimizing reprojection errors onto RGB frames, while Fixed-GS enables separate optimization of scene structure and color.
@@ 2 Related Works
Joint Pose and Scene Optimization. The research community has recently focused on developing methods sparf ; localrf ; nerfmm ; graf ; nopenerf ; barf ; cf3dgs that can be optimized without requiring precomputed camera poses. A line of work has focused on improving the stability of the optimization process. GARF graf and BARF barf both find that the high-frequency position encoding is prone to local minima and try to improve it. For example, GARF graf proposes using Gaussian activation to replace the sinusoidal position encoding. Another line of work has investigated incorporating additional constraints to make the problem more tractable. LocalRF localrf leverages the prior assumption of continuous motion between adjacent frames and progressively adds and optimizes camera poses. More recent approaches nopenerf ; localrf ; cf3dgs leverage pre-trained networks, i.e., monocular depth estimation and optical flow estimation. Exploiting 3DGS’s explicit representation, CF-3DGS cf3dgs directly back-projects Gaussian points using depth maps. While the aforementioned methods have made notable progress, they have yet to fully address the challenges posed by high-speed scenarios or rely on a good pose initialization. Our approach addresses these issues by leveraging motion and brightness cues from event streams.
@@
Event-Based Novel View Synthesis. Recent works have explored the integration of event cameras EMoTive ; etap ; urnet into the NeRF or 3DGS framework. Early approaches, such as E-NeRF enerf and EventNeRF eventnerf , utilize event-based generative models, minimizing the difference between the rendered brightness changes and observed brightness changes. Building upon this, Robust e-NeRF robustenerf incorporates a more realistic imaging model into the event-based framework, accounting for factors like refractory periods and noise. Beyond event-based NeRF, efforts have also been made to integrate event data into image-based methods. For instance, E2NeRF E2NeRF and EvDeblurNeRF evdeblurnerf leverage the Event Double Integral (EDI) edi model to address the deblurring problem, while DE-NeRF denerf and EvDNeRF EvDNeRF leverage the high temporal resolution property of event cameras to capture fast-moving elements in dynamic scene. More recently, Event-3DGS event3dgs and EaDeblur-GS eadeblurgs have extended previous approaches to 3D Gaussian Splatting, achieving superior rendering quality and real-time performance. A key distinction of our work is that, unlike the prior methods that rely on accurate precomputed poses, we target free-trajectory scenarios, jointly optimizing for both the camera poses and the scene representation. Furthermore, while previous works have been limited to simulated and simple environments, we evaluate our approach in large-scale outdoor scenarios with complex motions and lighting conditions.
@@
Figure 2: Method overview. The inputs are video frames and event stream. In the first stage, we progressively add new event images, leveraging the events and most recent frame to establish the event-driven optimization. In the second stage, we adopt the Fixed-GS strategy to mitigate the color distortion of 3DGS. The details of $\mathcal{L}_{LEGM}$ and CMax framework are shown in Fig. 3.
@@ 3 Preliminary
3DGS 3dgs parametrizes the 3D scene as a set of 3D gaussians $\{G_{k}\}_{k=1}^{K}$ that carry the geometric and appearance information. Each 3D Gaussian is characterized by several learnable properties, including its center position $\mu\in\mathbb{R}^{3}$, opacity $\alpha\in[0,1]$, spherical harmonics (SH) features $\mathbf{f}_{k}\in\mathbb{R}^{3\times 16}$ for view-dependent color $c\in\mathbb{R}^{3}$, rotation matrix $R\in\mathbb{R}^{3\times 3}$ (stored in quaternion form), scale factor $s\in\mathbb{R}^{3}$. The shape of each Gaussian is defined by the covariance matrix $\mathbf{\Sigma}$ and the center (mean) point $\mu$, $G(x)=\exp(-\frac{1}{2}{(x-\mu)}^{T}\mathbf{\Sigma}^{-1}(x-\mu)).$ During rendering, a tile-based rasterizer is applied to enable fast sorting and $\alpha$-blending. The color of each pixel is calculated via blending N ordered overlapping points:
@@
where $c_{i}$ is calculated from spherical harmonics and view direction, $\alpha_{i}$ is the multiplication of opacity and the transformed 2D Gaussian and r denotes the image pixel. With the forward rendering procedure, we can optimize 3DGS by minimizing a weighted combination loss of $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ between observation and rendered pixels: $\mathcal{L}_{color}=(1-\lambda)\mathcal{L}_{1}(\hat{I},I)+\lambda\mathcal{L}_{D-SSIM}(\hat{I},I),$ where $\lambda$ is balancing weight which is set to 0.2 following 3dgs . By integrating depth $d_{i}$ in Equation˜1 along the ray, we can also obtain a expected depth value $\hat{D}(\mathbf{r})$:
@@ 4 Method
[…]
@@ 4.1 EGM Driven Optimization
The EGM describes how event cameras asynchronously record pixel-wise brightness changes. When the logarithmic brightness change at a pixel $\mathbf{u}_{k}=(x_{k},y_{k})$, exceeds a predefined contrast threshold $C$,
@@
Figure 3: The illustration of unified CMax and LEGM optimization. We warp previous event frames to the sampled timestamp through the optical flow and maximize the sharpness of the image of IPWE. The byproduct IPWE is utilized to establish additional constraints on 3DGS.
@@ 4.2 Unified CMax and LEGM Optimization
While $\mathcal{L}_{EGM}$ leverages the brightness change information recorded by events, it does not explicitly exploit the motion information encoded in the event stream. To address this, we introduce the Contrast Maximization (CMax) cmax ; cmax2 ; cmax3 ; cmaxslam framework and the LEGM legm2015 ; legm2018 ; linearinversed . These models complement the previous EGM-driven optimization.
Under constant scene illumination, events are triggered by the motion of scene edges, forming continuous trajectories in (x, y, t) space. As shown in Fig. 2 (1.2), by warping(back-projecting) events along the correct motion trajectories, we can obtain a sharp image of warped events (IWE). Therefore, the sharpness of the IWE can serve as an indication of the accuracy of the estimated motion. This insight motivates us to derive the motion field by leveraging the rendered depth from 3DGS using eq.˜2 and the relative camera motion between neighboring timestamps. By optimizing the sharpness of the IWE, we can obtain the optimal motion field, which in turn helps to improve the geometric accuracy of the 3DGS and the camera poses.
@@
$E_{i,j-m\rightarrow j}=\textrm{warp}(E_{i,j-m},F_{i,j\rightarrow j-m}),$ (7)
where $m\in[0,r]$, $F_{i,j-m\rightarrow j}$ is the optical flow derived from the rendered depth $\hat{D}$ in  eq.˜2 and relative pose $T_{i,j\rightarrow j-m}$ between two timestamps:
$F_{i,j\rightarrow j-m}=\Pi(T_{i,j\rightarrow j-m}\Pi^{-1}(x,y,\hat{D}))-(x,y),$ (8)
@@
Furthermore, based on the LEGM legm2015 ; legm2018 , the brightness change $\Delta L$ at pixel $\mathbf{u}$ can be approximated by the dot product of the image gradient $\nabla L$ and the optical flow $\dot{\mathbf{u}}$ (note that $L$ is the logarithm of an image):
@@
It is noteworthy that the IPWE also encodes brightness change information. Combining  eq.˜9 and  eq.˜11, we establish a connection between the IPWE and the brightness changes of the rendered images:
@@
Note that to compute $F_{i,j\rightarrow j+1}$, we estimate $T_{i,j+1}$ by leveraging the assumption of locally linear motion from $T_{i,j-1}$ and $T_{i,j}$. Based on this relationship, we formulate an additional gradient-based loss:
[…]